OpenStudy (anonymous):

If w(t) = 3t + 5, solve w(t - 1) = w(2t)

OpenStudy (anonymous):

i think u subsitute 3t+5 on w and then solve like any equation

OpenStudy (lalaly):

w(t-1) = 3(t-1)+5= 3t-3+5 = 3t+2 w(2t) = 3(2t)+5= 6t+6 now w(t-1)=w(2t) 3t+2= 6t+5 -3t=3 t=-1

OpenStudy (anonymous):

what did u do? i didnt understand

OpenStudy (lalaly):

w(t-1) u substitute t-1 instead of t in w(t) and the same fr w(2t)

OpenStudy (anonymous):

ohh ok thanks!